# UOE: Unlearning one Expert is Enough for Mixture-of-Experts LLMs

### v1

@@ Abstract
Recent advancements in large language model (LLM) unlearning have shown remarkable success in removing unwanted data-model influences while preserving the model’s utility for legitimate knowledge. However, despite these strides, sparse Mixture-of-Experts (MoE) LLMs–a key subset of the LLM family–have received little attention and remain largely unexplored in the context of unlearning. As MoE LLMs are celebrated for their exceptional performance and highly efficient inference processes, we ask: How can unlearning be performed effectively and efficiently on MoE LLMs? And will traditional unlearning methods be applicable to MoE architectures? Our pilot study shows that the dynamic routing nature of MoE LLMs introduces unique challenges, leading to substantial utility drops when existing unlearning methods are applied. Specifically, unlearning disrupts the router’s expert selection, causing significant selection shift from the most unlearning target-related experts to irrelevant ones. As a result, more experts than necessary are affected, leading to excessive forgetting and loss of control over which knowledge is erased. To address this, we propose a novel single-expert unlearning framework, referred to as UOE, for MoE LLMs. Through expert attribution, unlearning is concentrated on the most actively engaged expert for the specified knowledge. Concurrently, an anchor loss is applied to the router to stabilize the active state of this targeted expert, ensuring focused and controlled unlearning that preserves model utility. The proposed UOE framework is also compatible with various unlearning algorithms. Extensive experiments demonstrate that UOE enhances both forget quality up to $5\%$ and model utility by $35\%$ on MoE LLMs across various benchmarks, LLM architectures, while only unlearning $0.06\%$ of the model parameters.
@@ 1 Introduction
To the best of our knowledge, the problem (Q) remains unexplored in the current literature. Our investigation begins with a pilot study that applies existing unlearning methods to MoE LLMs. Preliminary results indicate that a simple implementation of these methods can lead to a substantial drop in model utility and even model collapse. This phenomenon is illustrated in Fig. 1(a), which depicts the performance of the unlearned MoE LLMs predominantly closer to the bottom right corner, indicating a significant and unacceptable utility drop compared to conventional dense LLMs.
To fully understand this phenomenon, we begin by performing a careful sanity check on unlearning methods in MoE LLMs and conduct an in-depth analysis of failure cases. Ideally, in MoE LLMs, given an input, the routers should evaluate and direct it to the most relevant experts, with unlearning targeting and erasing the corresponding knowledge in these experts. However, by monitoring expert selection during unlearning, we find that the process often prompts routers to constantly switch the activated experts. This behavior persists even when routers are fixed. As a result, unlearning algorithms create “short-cuts”, where instead of targeting the most relevant experts, the routers shift to less relevant ones to trick for unlearning loss reduction (i.e., irrelevant experts could be unlearned). This leads to substantial drops in model utility. See Fig. 1(b) for illustration.
@@
Figure 1: Overview of the key findings in this paper. (a) Illustration of the ineffectiveness of existing unlearning methods on MoE LLMs. Four unlearning algorithms—GA (Eldan & Russinovich, 2023), GDiff (Maini et al., 2024), NPO (Zhang et al., 2024), and RMU (Li et al., 2024)—were applied to two MoE LLMs (DeepSeek-v2-Lite (Liu et al., 2024a) and Qwen1.5-MoE (Team, 2024)) and two dense LLMs (Phi3.5 (Abdin et al., 2024) and LLaMA3-8B (Dubey et al., 2024)) using the WMDP benchmark (Li et al., 2024). The drop in target knowledge (accuracy drop on the forget test set, higher is better) and the drop in model utility (accuracy drop on MMLU (Hendrycks et al., 2023), lower is better) are plotted. Ideal performance is in the top left corner, but MoE LLMs show poor unlearning quality with sharp utility drop. (b) Illustration of ideal versus ineffective MoE LLM unlearning. Target experts—those most frequently activated given the forget set—are identified for unlearning. However, existing unlearning algorithms tend to cause substantial expert selection shifts, leading to excessive and unnecessary unlearning of non-target experts, which significantly impairs model utility.
To solve the problem, we propose a novel unlearning framework specifically tailored for MoE LLMs, named UOE, which stands for Unlearning One Experts. UOE employs expert attribution to pinpoint the expert most actively involved with the forget set, which is designated as the primary target for unlearning. Unlearning efforts are exclusively focused on this identified expert. Concurrently, an anchor loss is applied to the router to stabilize the active status of the targeted expert throughout the unlearning process. This approach prevents the frequent switching of expert selection, ensuring that unlearning is both focused and controlled. Our contributions are summarized below.
@@
$\bullet$ We propose a novel parameter-efficient unlearning framework, UOE, tailored for MoE LLMs. UOE effectively pinpoints, fixates, and unlearns the most pertinent expert relative to the forget set. UOE enjoys high flexibility and works in a plug-in-and-play mode with any existing unlearning methods to boost forget quality, model utility, and efficiency at the same time.
$\bullet$ We conduct extensive experiments to demonstrate the effectiveness of UOE across various MoE architectures, MU benchmarks, and unlearning methods. Our results show that when integrated with UOE, all tested unlearning methods achieve significant improvements in model utility up to $35\%$ and concurrently enhance the quality of forgetting with only $0.06\%$ parameters being updated.
@@
Machine Unlearning for LLMs. A growing body of research has investigated the problem of unlearning in large language models (LLMs) (Yao et al., 2024; Lu et al., 2022; Jang et al., 2022; Kumar et al., 2022; Zhang et al., 2023a; Pawelczyk et al., 2023; Eldan & Russinovich, 2023; Ishibashi & Shimodaira, 2023; Yao et al., 2023; Maini et al., 2024; Zhang et al., 2024; Li et al., 2024; Wang et al., 2024a; Jia et al., 2024b; Liu et al., 2024c; b; Thaker et al., 2024). These studies have practical applications, such as removing sensitive information (Jang et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023) and preventing the generation of harmful or biased content (Jang et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023; Lu et al., 2022; Yu et al., 2023; Yao et al., 2023; Liu et al., 2024d), memorized sequences (Jang et al., 2022; Barbulescu & Triantafillou, 2024), and copyrighted material (Eldan & Russinovich, 2023; Jang et al., 2022). To facilitate unlearning, recent methods aim to bypass the need for retraining models from scratch by excluding the forget set containing the targeted data to be removed (Ilharco et al., 2022; Liu et al., 2022; Yao et al., 2023; Eldan & Russinovich, 2023; Jia et al., 2024b; Zhang et al., 2024; Li et al., 2024; Thaker et al., 2024; Liu et al., 2024b). Techniques like task arithmetic also enable efficient model editing through parameter merging (Hu et al., 2024; Ilharco et al., 2022). Although these methods do not provide exact unlearning akin to full retraining, they remain efficient and effective under empirical unlearning evaluation metrics. Approaches often include model fine-tuning and optimization (Liu et al., 2022; Yao et al., 2023; Eldan & Russinovich, 2023; Jia et al., 2024b; Zhang et al., 2024; Li et al., 2024), or input prompting and in-context learning (Thaker et al., 2024; Pawelczyk et al., 2023; Liu et al., 2024b). Other approaches, such as localization-informed unlearning, identify and locally edit model units (e.g., layers or neurons) closely related to the data or tasks being unlearned (Meng et al., 2022; Wu et al., 2023; Wei et al., 2024b). Despite these efforts, studies have shown that forgotten knowledge can often still be extracted from models post-unlearning (Patil et al., 2024; Liu et al., 2024e; Lynch et al., 2024; Shostack, 2024). However, most existing research has focused on dense LLMs, leaving unlearning in MoE LLMs largely unexplored. For example, the unlearning of Mixtral-$8\times 7B$ is discussed in Li et al. (2024), but only a single method with ad-hoc adjustments was examined. This work aims to fill this gap by conducting a comprehensive study of various unlearning methods, benchmarks, and MoE models, addressing the specific challenges posed by the MoE architecture.
@@
MoE-based LLMs. Sparse Mixture-of-Experts (MoE) models are designed to activate only a subset of expert networks for each input, enabling substantial model scaling with minimal computational overhead (Shazeer et al., 2017). Current approaches to MoE model development can be categorized into two types: training from scratch (Fedus et al., 2022; Zoph et al., 2022a; Shen et al., 2023) and building from dense checkpoints (Zhang et al., 2021; Komatsuzaki et al., 2022; Zhu et al., 2024). Over recent years, MoE models have seen key advancements, including improvements in scalability (Riquelme et al., 2021; Kim et al., 2021; Zhou et al., 2022; Zoph et al., 2022a), efficiency optimization (Fedus et al., 2022; Lepikhin et al., 2020; Chowdhery et al., 2023), and expert balancing techniques (Cong et al., 2024; Zoph et al., 2022b; Dai et al., 2022). The implementation of transformer-based MoE models has been successfully integrated into LLMs, significantly enhancing inference efficiency (Jiang et al., 2024; Dai et al., 2024; Team, 2024; xAI, 2024; Hong et al., 2024; Abdin et al., 2024; Lieber et al., 2024; Yang et al., 2024; Zhu et al., 2024; Databricks, 2024; Xue et al., 2024). For example, DeepSeekMoE (Dai et al., 2024) improves expert specialization by segmenting experts into smaller subsets for flexible activation, while isolating shared experts to reduce redundancy and capture common knowledge. Similarly, Qwen1.5-MoE (Team, 2024) partitions a standard FFN layer into smaller segments to create multiple experts, introducing a fine-grained routing mechanism that enables Qwen1.5-MoE to match the performance of 7B models while using only one-third of the activation parameters. Despite the efficiency gains provided by MoE’s dynamic routing system, existing research highlights additional challenges compared to traditional dense models, including unstable training (Zoph et al., 2022a; Dai et al., 2022), robustness issues (Zhang et al., 2023b; Puigcerver et al., 2022), and complications in parallel deployment (Hwang et al., 2023; Gale et al., 2023). In this work, we show that the root cause of the ineffectiveness of existing unlearning methods for MoE LLMs also stems from the dynamic routing system.
@@ 3 Preliminaries
In this section, we start by presenting the mathematical formulation of LLM unlearning. The lack of exploration on MoE LLM unlearning inspires us to investigate whether existing unlearning methods keep effective in these models. Our pilot study reveals that methods designed for conventional LLMs are ineffective in unlearning MoE LLMs.
Preliminaries on MoE LLM unlearning. Based on the generic formulation of LLM unlearning outlined in Liu et al. (2024c), the task of LLM unlearning can be formulated as eliminating the influence of a specific ‘unlearning target’–whether it is related to data, knowledge, or model capabilities–from a pretrained LLM (denoted by $\bm{\theta}_{o}$). The unlearning target is typically defined by a forget set $\mathcal{D}_{f}$, which contains the information or knowledge to be removed. To ensure the model retains its generation ability (i.e., utility) after unlearning, a retain set $\mathcal{D}_{r}$ is introduced, consisting of data unrelated to the unlearning target. With this setup, the LLM unlearning problem is usually formed as a regularized optimization problem, finetuned from $\bm{\theta}_{o}$ using both the forget set $\mathcal{D}_{f}$ and the retain set $\mathcal{D}_{r}$:
@@
Next, we provide a brief introduction to how the routing system operates in the MoE LLM architecture. In MoE LLMs, e.g., DeepSeek-v2-Lite (Liu et al., 2024a), the feed-forward networks (FFNs) of Transformers are split into multiple experts and activated by the output of the router in front of the expert layers, see Fig. 1(b) for illustration. In the $l$-th layer, given the input $\mathbf{u}_{t}^{(l)}$ corresponding to the $t$-th token, router layers calculate the score of each token and assign them to top-$K$ experts:
@@
Table 1: Unlearning performance when controlling tunable parameters in MoE LLMs.
Tunable Module Forget Quality $\downarrow$ Retain Quality $\uparrow$
Qwen
Original 0.4192 0.5979
@@
Unlearning for MoE LLM is not trivial: a pilot study. The goal of unlearning is twofold: (1) to ensure the model forgets the targeted information and knowledge stored in $\mathcal{D}_{f}$, and (2) to preserve the model utility without significant degradation. Our pilot study reveals that the special routing system in MoE LLMs introduces additional challenges to unlearning, rendering existing methods ineffective. We applied four widely used LLM unlearning methods: GA (Gradient Ascent) (Eldan & Russinovich, 2023), GDiff (Gradient Difference) (Maini et al., 2024), NPO (Negative Preference Optimization) (Zhang et al., 2024), and RMU (Representation Misdirection for Unlearning) (Li et al., 2024) with the WMDP benchmark (Li et al., 2024) on two MoE LLMs, Qwen1.5-MoE (Team, 2024) and DeepSeek-V2-Lite (Liu et al., 2024a), as well as two dense LLMs for reference, LLaMA3-8B (Dubey et al., 2024) and Phi-3.5-mini-instruct (Abdin et al., 2024), where the task aims to unlearn hazardous knowledge in LLMs. In Fig. 1(a), to ease the comparison, we report the forget quality (performance drop on the forget test set, where higher is better) against retain quality (performance drop on the MMLU (Hendrycks et al., 2020) utility benchmark, where lower is better). Each data point represents the best result of a model-method combination with hyper-parameter tuning, with ideal performance located near the top left corner, signifying high unlearning effectiveness with minimal impact on model utility. As we can see, most MoE LLM data points cluster in the lower right, indicating severe utility drops and poor unlearning performance compared to dense models. In Fig. 1(a), all model parameters (including routers and experts) are involved in unlearning. To ensure that these poor results are not due to improper parameter settings, Tab. 1 presents additional experiments using two other parameter configurations (routers-only and experts-only) for GA, yet no significant improvements are observed in either forget or retain quality (more than 20% utility drop). The results above imply the problem of MoE LLM unlearning is more challenging and far from trivial, even if LLM unlearning is well-studied.
4 Our Proposal: Unlearning one Expert (UOE)
In this section, we delve into the failure cases highlighted in Sec. 3 by analyzing the behavior of routers and their expert selection patterns. We then identify two primary root causes underlying the poor unlearning performance in MoE LLMs. Based on these insights, we introduce UOE, a new unlearning paradigm designed to achieve controllable and effective unlearning for MoE LLMs.

### v2

# SEUF: Is Unlearning One Expert Enough for Mixture-of-Experts LLMs?

@@ Abstract
Recent advancements in LLMs unlearning have shown remarkable success in removing unwanted data-model influences while preserving the model’s utility for legitimate knowledge. Despite these strides, sparse Mixture-of-Experts (MoE) LLMs–a key subset of the LLM family–have remain unexplored in the context of unlearning. As MoE LLMs are celebrated for their exceptional performance, we ask: How can unlearning be performed effectively and efficiently on MoE LLMs? Our pilot study shows that the dynamic routing nature of MoE LLMs introduces unique challenges, leading to excessive forgetting, uncontrolled knowledge erasure and substantial utility drops when existing unlearning methods are applied. To address this, we propose a novel Selected-Expert Unlearning Framework (SEUF). Through expert attribution, unlearning is concentrated on the most actively engaged experts for the specified knowledge. Concurrently, an anchor loss is applied to the router to stabilize the active state of this targeted expert, ensuring focused and controlled unlearning. SEUF is compatible with various standard unlearning algorithms. Extensive experiments demonstrate that SEUF enhances both forget quality up to $5\%$ and model utility by $35\%$ on MoE LLMs across various benchmarks and LLM architectures (compared to standard unlearning algorithms), while only unlearning $0.06\%$ of the model parameters.
SEUF: Is Unlearning One Expert Enough for Mixture-of-Experts LLMs?
Haomin Zhuang1∗, Yihua Zhang2thanks: Equal contribution, Kehan Guo1, Jinghan Jia2, Gaowen Liu3, Sijia Liu2, Xiangliang Zhang1 1University of Notre Dame 2Michigan State University 3Cisco Research {hzhuang2,xzhang33}@nd.edu
@@ 1 Introduction
To the best of our knowledge, the problem (Q) remains unexplored in the current literature. Our investigation begins with a pilot study that applies existing unlearning methods to MoE LLMs. Preliminary results indicate that a simple application of these methods can lead to a substantial drop in model utility and even model collapse. This phenomenon is illustrated in Fig. 1(a), which depicts the performance of the unlearned MoE LLMs predominantly closer to the bottom right corner, indicating a significant and unacceptable utility drop compared to conventional dense LLMs.
To look into this phenomenon, we begin by performing a careful sanity check on unlearning methods in MoE LLMs and conduct an in-depth analysis of failure cases. Ideally, in MoE LLMs, given an input, the routers should evaluate and direct it to the most relevant experts, with unlearning targeting and erasing the corresponding knowledge in these experts. However, by monitoring expert selection during unlearning, we find that the process often prompts routers to constantly switch the activated experts. This behavior persists even when routers are fixed. As a result, unlearning algorithms create “short-cuts”, where instead of targeting the most relevant experts, the routers shift to less relevant ones to trick for unlearning loss reduction (i.e., irrelevant experts are unlearned). This leads to substantial drops in model utility (illustrated in Fig. 1(b)).
@@
Figure 1: Overview of the key findings in this paper. (a) Illustration of the ineffectiveness of existing unlearning methods on MoE LLMs. Four unlearning algorithms—GA (Eldan and Russinovich, 2023), GDiff (Maini et al., 2024), NPO (Zhang et al., 2024), and RMU (Li et al., 2024)—were applied to two MoE LLMs (DeepSeek-v2-Lite (Liu et al., 2024a) and Qwen1.5-MoE (Team, 2024)) and two dense LLMs (Phi3.5 (Abdin et al., 2024) and LLaMA3-8B (Dubey et al., 2024)) using the WMDP benchmark Li et al. (2024). The drop in target knowledge (accuracy drop on the forget test set, higher is better) and the drop in model utility (accuracy drop on MMLU Hendrycks et al. (2023), lower is better) are plotted. Better-unlearned models should appear in the top left corner, but unlearning on MoE LLMs was less effective compared to non-MoE modles. (b) Illustration of ideal versus ineffective MoE LLM unlearning. Target experts—those most frequently activated given the forget set—are identified for unlearning. However, existing unlearning algorithms tend to cause substantial expert selection shifts, leading to excessive and unnecessary unlearning of non-target experts, which significantly impairs model utility.
To solve the problem, we propose a novel unlearning framework specifically tailored for MoE LLMs, named SEUF (Selected Experts Unlearning Framework). SEUF employs expert attribution to pinpoint the experts most actively involved with the forget set, which is designated as the primary target for unlearning. Unlearning efforts are exclusively focused on this identified expert. Concurrently, an anchor loss is applied to the router to stabilize the active status of the targeted expert throughout the unlearning process. This approach prevents the frequent switching of expert selection, ensuring that unlearning is both focused and controlled. Our contributions are summarized below.
@@
$\bullet$ We propose a novel parameter-efficient unlearning framework, SEUF, for MoE LLMs. SEUF effectively pinpoints, fixates, and unlearns the most pertinent experts relative to the forget set. SEUF enjoys high flexibility and works in a plug-in-and-play mode with any existing unlearning methods to boost forget quality, model utility, and efficiency at the same time.
$\bullet$ We conduct extensive experiments to demonstrate the effectiveness of SEUF across various MoE architectures, MU benchmarks, and unlearning methods. Our results show that when integrated with SEUF, all tested unlearning methods achieve substantial improvements in model utility up to $35\%$ and concurrently enhance the quality of forgetting with only $0.06\%$ parameters being updated.
@@ 2 Related Works
Machine Unlearning for LLMs. A growing body of research has investigated the problem of unlearning in LLMs (Yao et al., 2024; Lu et al., 2022; Jang et al., 2022; Kumar et al., 2022; Zhang et al., 2023a; Pawelczyk et al., 2023; Eldan and Russinovich, 2023; Ishibashi and Shimodaira, 2023; Yao et al., 2023; Maini et al., 2024; Zhang et al., 2024; Li et al., 2024; Wang et al., 2024a; Jia et al., 2024b; Liu et al., 2024c, b; Thaker et al., 2024). These studies have practical applications, such as removing sensitive information (Jang et al., 2022; Eldan and Russinovich, 2023; Wu et al., 2023), preventing the generation of harmful or biased content (Jang et al., 2022; Eldan and Russinovich, 2023; Wu et al., 2023; Lu et al., 2022; Yu et al., 2023; Yao et al., 2023; Liu et al., 2024d), memorized sequences (Jang et al., 2022; Barbulescu and Triantafillou, 2024), and copyrighted material (Eldan and Russinovich, 2023; Jang et al., 2022). To facilitate unlearning, recent methods aim to bypass the need for retraining models from scratch by excluding the forget set containing the targeted data to be removed (Ilharco et al., 2022; Liu et al., 2022; Yao et al., 2023; Eldan and Russinovich, 2023; Jia et al., 2024b; Zhang et al., 2024; Li et al., 2024; Thaker et al., 2024; Liu et al., 2024b). Techniques like task arithmetic also enable efficient model editing through parameter merging (Hu et al., 2024; Ilharco et al., 2022). Although these methods do not provide exact unlearning akin to full retraining, they remain efficient and effective under empirical unlearning evaluation metrics. Approaches often include model fine-tuning and optimization (Liu et al., 2022; Yao et al., 2023; Eldan and Russinovich, 2023; Jia et al., 2024b; Zhang et al., 2024; Li et al., 2024), or input prompting and in-context learning (Thaker et al., 2024; Pawelczyk et al., 2023; Liu et al., 2024b). Other approaches, such as localization-informed unlearning, identify and locally edit model units (e.g., layers or neurons) closely related to the data or tasks being unlearned (Meng et al., 2022; Wu et al., 2023; Wei et al., 2024). Most existing research has focused on dense LLMs, leaving unlearning in MoE LLMs unexplored. For example, the unlearning of Mixtral-$8\times 7B$ discussed in Li et al. (2024) only examined a single method with ad-hoc adjustments. This work aims to fill this gap by conducting a comprehensive study of various unlearning methods, benchmarks, and MoE models, addressing the specific challenges posed by the MoE architecture.
@@
MoE-based LLMs. Sparse MoE models are designed to activate only a subset of expert networks for each input during inference, enabling substantial model scaling with minimal computational overhead (Shazeer et al., 2017). Current MoE model development can be categorized into two types: training from scratch (Fedus et al., 2022; Zoph et al., 2022a; Shen et al., 2023) and building from dense checkpoints (Zhang et al., 2021; Komatsuzaki et al., 2022; Zhu et al., 2024). Over recent years, MoE models have seen key advancements, including improvements in scalability (Riquelme et al., 2021; Kim et al., 2021; Zhou et al., 2022; Zoph et al., 2022a), efficiency optimization (Fedus et al., 2022; Lepikhin et al., 2020; Chowdhery et al., 2023), and expert balancing techniques (Cong et al., 2024; Zoph et al., 2022b; Dai et al., 2022). The implementation of transformer-based MoE models has been integrated into LLMs, significantly enhancing inference efficiency (Jiang et al., 2024; Dai et al., 2024; xAI, 2024; Hong et al., 2024; Abdin et al., 2024; Lieber et al., 2024; Yang et al., 2024; Zhu et al., 2024; Databricks, 2024). For example, DeepSeekMoE (Dai et al., 2024) improves expert specialization by segmenting experts into smaller subsets for flexible activation, while isolating shared experts to reduce redundancy and capture common knowledge. Similarly, Qwen1.5-MoE (Team, 2024) partitions a standard FFN layer into smaller segments to create multiple experts, introducing a fine-grained routing mechanism that enables Qwen1.5-MoE to match the performance of 7B models with only one-third of parameters activated. Despite the efficiency gains provided by MoE’s dynamic routing system, existing research highlights additional challenges compared to traditional dense models, including unstable training (Zoph et al., 2022a; Dai et al., 2022), robustness issues (Zhang et al., 2023b; Puigcerver et al., 2022), and complications in parallel deployment (Hwang et al., 2023; Gale et al., 2023). In this work, we show that the root cause of the ineffectiveness of existing unlearning methods for MoE LLMs also stems from the dynamic routing system.
@@ 3 Preliminaries
In this section, we present our pilot study to reveal that unlearning methods designed for conventional LLMs are ineffective in unlearning MoE LLMs.
Preliminaries on MoE LLM unlearning. Based on the generic formulation outlined in Liu et al. (2024c), the task of LLM unlearning is to eliminate the influence of a specific ‘unlearning target’–whether it is related to data, knowledge, or model capabilities–from a pretrained LLM (denoted by $\bm{\theta}_{o}$). The unlearning target is typically defined by a forget set $\mathcal{D}_{f}$, which contains the information or knowledge to be removed. To ensure the model retains its generation ability (i.e., utility) after unlearning, a retain set $\mathcal{D}_{r}$ is introduced, consisting of data unrelated to the unlearning target. With this setup, the LLM unlearning problem is usually formed as a regularized optimization problem, finetuned from $\bm{\theta}_{o}$ using both the forget set $\mathcal{D}_{f}$ and the retain set $\mathcal{D}_{r}$:
@@
Next, we provide a brief introduction to how the routing system operates in the MoE LLM architecture. In MoE LLMs, e.g., DeepSeek-v2-Lite (Liu et al., 2024a), the feed-forward networks (FFNs) of Transformers are split into multiple experts and activated by the output of the router in front of the expert layers, see Fig. 1(b) for illustration. In the $l$-th layer, given the input $\mathbf{u}_{t}^{(l)}$ corresponding to the $t$-th token, router layers calculate the score of each token and assign them to the top-$K$ experts:
@@
Table 1: Unlearning performance of GA when controlling tunable parameters in MoE LLMs.
Tunable Module Forget Efficacy $\downarrow$ Utility $\uparrow$
Qwen Original 0.4192 0.5979
@@
Unlearning for MoE LLM is not trivial: a pilot study. The goal of unlearning is twofold: (1) to ensure the model forgets the targeted information and knowledge stored in $\mathcal{D}_{f}$, and (2) to preserve the model utility without significant degradation. Our pilot study reveals that the special routing system in MoE LLMs introduces additional challenges to unlearning, rendering existing methods ineffective. We applied four widely used LLM unlearning methods: GA (Gradient Ascent) (Eldan and Russinovich, 2023), GDiff (Gradient Difference) (Maini et al., 2024), NPO (Negative Preference Optimization) (Zhang et al., 2024), and RMU (Representation Misdirection for Unlearning) (Li et al., 2024) with the WMDP benchmark (Li et al., 2024) on two MoE LLMs, Qwen1.5-MoE (Team, 2024) and DeepSeek-V2-Lite (Liu et al., 2024a), as well as two dense LLMs for reference, LLaMA3-8B (Dubey et al., 2024) and Phi-3.5-mini-instruct (Abdin et al., 2024), where the task aims to unlearn hazardous knowledge in LLMs. In Fig. 1(a), to ease the comparison, we report the forget quality (performance drop on the forget test set, where higher is better) against retain quality (performance drop on the MMLU (Hendrycks et al., 2020) utility benchmark, where lower is better). Each data point represents the best result of a model-method combination with hyper-parameter tuning, with ideal performance located near the top left corner, signifying high unlearning effectiveness with minimal impact on model utility. As we can see, most MoE LLM data points cluster in the lower right, indicating severe utility drops and poor unlearning performance compared to dense models. In Fig. 1(a), all model parameters (including routers and experts) are involved in unlearning. To ensure that these poor results are not due to improper parameter settings, Tab. 1 presents additional experiments using two other parameter configurations (routers-only and experts-only) for GA, yet no significant improvements are observed in either forget or retain quality (more than 20% utility drop). The results above imply the problem of MoE LLM unlearning is more challenging and far from trivial, even if LLM unlearning is well-studied.
4 Our Proposal: SEUF
In this section, we delve into the failure cases highlighted in Sec. 3 by analyzing the behavior of routers and their expert selection patterns. We then identify two primary root causes underlying the poor unlearning performance in MoE LLMs. Based on these insights, we introduce SEUF, a new unlearning paradigm designed to achieve controllable and effective unlearning for MoE LLMs.
Uncovering the root cause: ‘short-cut’ in MoE LLM unlearning and expert selection shift. In order to fully understand the failure cases of MoE LLM unlearning, we begin by inspecting and monitoring the expert selection pattern of the unlearned model. In Fig. 2, we show the proportion of tokens assigned to each selected expert on the data samples from WMDP forget dataset (Li et al., 2024). For the input of a specific topic, a small portion of experts (around 6 to 9 out of 64 experts) were assigned with the majority of the tokens in each layer, which was also confirmed in Wang et al. (2024b). Thus, we have the following insight: